OpenStudy (rissa):

Find the slope and y-intercept of the line. 3/2x - 5

OpenStudy (saifoo.khan):

y = mx + b where m is the slope and b is the intercept. so slope = 3/2 y-int = -5

OpenStudy (rissa):

i don't know why i keep forgetting the formula thank you!